### v2

# Simple yet Effective Semi-supervised Knowledge Distillation from Vision-Language Models via Dual-Head Optimization

Abstract
Semi-supervised learning (SSL) has emerged as a practical solution for addressing data scarcity challenges by leveraging unlabeled data. Recently, vision-language models (VLMs), pre-trained on massive image-text pairs, have demonstrated remarkable zero-/few-shot performance that often surpasses SSL approaches due to their exceptional generalization capabilities. This gap motivates us to question: how can we effectively harness the powerful generalization capabilities of VLMs into task-specific models? Knowledge distillation (KD) offers a natural framework for transferring VLM capabilities, but we identify that it suffers from gradient conflicts between supervised and distillation losses. To address this challenge, we propose Dual-Head Optimization (DHO), which introduces dual prediction heads for each distinct signal. We observe that DHO resolves gradient conflicts, enabling improved feature learning compared to single-head KD baselines, with practical benefits of minimal computational overhead and test-time hyperparameter tuning without retraining. Extensive experiments across 15 datasets show that DHO consistently outperforms KD baselines, often outperforming teacher models with smaller student models. DHO also achieves new state-of-the-art performance on both in-distribution ImageNet semi-supervised learning and out-of-distribution generalization across ImageNet variants. We publicly release our code and model checkpoints to facilitate future research at https://github.com/erjui/DHO.
1 Introduction
Vision-language models (VLMs), which learn joint vision-language representations through large-scale pre-training, have shown remarkable zero-shot capabilities across diverse tasks (Radford et al., 2021; Jia et al., 2021). Building upon these strong foundational capabilities, recent work has explored various adaptation strategies, including parameter-efficient approaches such as linear probing (Li et al., 2022; Huang et al., 2024), lightweight adapters (Zhang et al., 2021; Gao et al., 2024), and prompt-based fine-tuning methods (Zhou et al., 2022b; a; Lafon et al., 2025), demonstrating the potential of VLMs for data-limited visual recognition tasks.
In parallel, semi-supervised learning (SSL) has emerged as a practical approach to address data scarcity by leveraging both labeled and unlabeled data (Sohn et al., 2020; Assran et al., 2021; Cai et al., 2022; Zheng et al., 2023). While these methods have shown success to leverage large amounts of unlabeled data, they often struggle to match the impressive zero- and few-shot capabilities of large pre-trained VLMs (Liu et al., 2023b). This discrepancy highlights a fundamental limitation: traditional semi-supervised methods, despite their theoretical appeal, remain suboptimal compared to the rich representations learned by foundation models through massive-scale pre-training.
VLMs excel at zero- and few-shot generalization but may lack the fine-grained discriminative power needed for specific tasks, while models trained on limited labeled data capture task-specific patterns but generalize poorly. This complementary nature motivates us to integrate generalist VLM knowledge with task-specific supervision in semi-supervised learning settings. Therefore, the challenge naturally arises: how can we effectively transfer the powerful capabilities of large VLMs to task-specific models in semi-supervised settings?
Figure 1: (Left): DHO consistently outperforms single-head baselines on 11 datasets under 16-shot semi-supervised setting. The improvements are evaluated in comparison to the second-best one. (Right): DHO achieves new SoTA on ImageNet in both 1% and 10% labeled data setting, with fewer parameters.
Knowledge distillation (KD; Hinton, 2015) emerges as a natural solution to the challenge, as it offeres an efficient framework to transfer knowledge from large VLMs to student models while simultaneously leveraging task-specific patterns from limited labeled data. However, existing VLM distillation methods primarily focus on general-purpose training (Yang et al., 2024a; Vasu et al., 2024; Udandarao et al., 2024; Yang et al., 2024c) or employ multi-stage pipelines with unsupervised pre-distillation (Vemulapalli et al., 2024; Wu et al., 2025), which require additional task-specific fine-tuning. Conventional single-stage KD methods, such as logit distillation (Hinton, 2015; Chen et al., 2020), offer a more direct approach, but we find them suboptimal in semi-supervised settings.
Through analysis, we identify that this stems from a fundamental problem: gradient conflicts between the supervised loss (from limited labeled data) and the distillation loss (from teacher predictions). This misalignment is particularly severe in semi-supervised settings, where the strong and consistent distillation signal from the teacher can overwhelm the weak and potentially noisy signal from scarce labeled data. Such gradient conflicts ( and in Fig.˜3) are well-documented to impair effective feature learning (Yu et al., 2020a; Liu et al., 2021; Chen & Er, 2025), preventing the model from finding an optimal balance between task-specific adaptation and general knowledge transfer.
To address the above issue, we propose a simple yet effective distillation framework, DHO (Dual-Head Optimization), which jointly leverages labeled samples and the probabilistic outputs of the teacher model. Specifically, it learns two distinct heads, each optimized with a separate loss: the supervised and the KD loss, respectively. Our analysis reveals that DHO mitigates gradient conflicts both in the classification head and the shared feature extractor ( in Fig.˜3), arising from the two different training signals and improves feature representations compared to baselines (Sections˜3.3 and 8). Our framework also enables controlling the relative influence of supervised and teacher predictions through linear combination of both head outputs, whose effectiveness is demonstrated in Fig.˜6. Additionally, DHO enables post-training hyperparameter adjustment by tuning the linear combination weights at inference (Fig.˜9), eliminating costly training-time hyperparameter search.
We conduct extensive experiments across 15 different datasets including ImageNet (Russakovsky et al., 2015). The experimental results demonstrate the consistent improvement of DHO over conventional KD methods across all evaluated datasets (Fig.˜1-Left). DHO sometimes even outperforms the zero/few-shot teacher models with smaller student models with the same labeled data, demonstrating effective task-specific enhancement beyond teacher capabilities (Figs.˜4 and 5). Furthermore, DHO achieves new state-of-the-art (SoTA) performance on ImageNet semi-supervised setting, improving accuracy by 3% and 0.1% with 1% and 10% labeled data, respectively, while using fewer parameters (Fig.˜1-Right). Notably, DHO can be seamlessly integrated with existing adaptation techniques with minimal computational overhead, achieving new SoTA on Out-of-Distribution (OOD) tasks across ImageNet distribution-shifted variants (Tab.˜2 and Section˜E.5).
Our contributions and findings are summarized as follows:
  • We firstly identify gradient conflict when integrating VLMs’ general knowledge with task-specific supervision from limited data. To address this, we propose Dual-Head Optimization (DHO), which optimizes the supervised and distillation objectives in separate heads.
  • DHO effectively resolves gradient conflicts both in the classification head and the shared feature extractor, leading to improved feature representations. Our framework enables flexible post-training adjustment of dual head output weights with minimal computational overhead.
  • In extensive experiments on 15 datasets, DHO consistently outperforms baselines and sometimes surpasses teacher model performance. It establishes new SoTA for in-distribution ImageNet (e.g., +3%/+0.1% at 1%/10% labels) with fewer parameters (76M/767M), and also achieves SoTA OOD generalization across ImageNet variants when integrated with adaptation methods.
2 Method
2.1 Preliminaries
We begin with preliminaries: a brief background on VLMs, the problem formulation for few-/low-shot learning, and single-head KD baselines. We defer related work—vision–language pretraining, data-limited adaptation of VLMs, knowledge distillation, and dual-head methods—to Appendix˜A.
Background on VLMs.
[…]
Problem Formulation.
We focus on transferring knowledge from VLMs to task-specific models under few-shot or low-shot semi-supervised learning scenarios, where both labeled and unlabeled data are utilized. Specifically, given a $K$-shot and $C$-class classification problem, we are provided with a labeled dataset $\mathcal{D}^{(l)}=\{(x^{(l)}_{n},y_{n})\}_{n=1}^{N}$, where $N=K\times C$ is the total number of labeled examples, and $y_{n}\in\{1,\ldots,C\}$ denotes the class labels. Additionally, we have access to an unlabeled dataset $\mathcal{D}^{(u)}=\{x^{(u)}_{m}\}_{m=1}^{M}$ consisting of $M$ unlabeled images. Low-shot learning represents a more realistic setting than traditional few-shot learning where only a small fraction, e.g., 1% ($\frac{N}{N+M}\approx 0.01$) or 10% ($\frac{N}{N+M}\approx 0.1$) of the total dataset is labeled. Our goal is then to develop a student model by leveraging $\mathcal{D}^{(l)}$ and $\mathcal{D}^{(u)}$, guided by the knowledge of the VLM encoders $f_{\mathcal{X}}$ and $f_{\mathcal{T}}$. The student model consists of a feature extractor $g(x)$ parameterzied by $\theta$ and a linear prediction head $h(z)=Wz+b\in\mathbb{R}^{c}$, where $W\in\mathbb{R}^{C\times d},b\in\mathbb{R}^{C}$, followed by the softmax function $\sigma$.
Figure 2: Conceptual illustration on KD frameworks, Single-Head Optimization (SHO) and Dual-Head Optimization (DHO), for semi-supervised settings. As demonstrated in Fig.˜3, we observe gradient conflict of SHO. In contrast, DHO mitigates such conflicts by leveraging dual-head architectures in Fig.˜3.
Single-head optimization (SHO) of KD.
Our method builds on logit distillation in semi-supervised settings (Hinton, 2015; Chen et al., 2020) that combines supervised loss $\mathcal{L}_{\text{CE}}$ on the labeled dataset $\mathcal{D}^{(l)}$ with KD loss $\mathcal{L}_{\text{KD}}$ on both labeled and unlabeled datasets $\mathcal{D}^{(l)}\cup\mathcal{D}^{(u)}$, i.e., $\lambda\mathcal{L}_{\text{CE}}+(1-\lambda)\mathcal{L}_{\text{KD}}$. Specifically, the supervised loss $\mathcal{L}_{\text{CE}}$ and the KD loss $\mathcal{L}_{\text{KD}}$ are defined as follows:
$\mathcal{L}_{\text{CE}}$ $=\frac{1}{N}\sum_{n}\ell\left(\sigma(h(z_{n}^{(l)})),y_{n}\right),$ (1)
$\mathcal{L}_{\text{KD}}$ $=\frac{1}{N}\sum_{n}D_{\mathrm{KL}}\left[p_{n}^{(l)}\|\sigma(h(z_{n}^{(l)}))\right]+\frac{1}{M}\sum_{m}D_{\mathrm{KL}}\left[p_{m}^{(u)}\|\sigma(h(z_{m}^{(u)}))\right].$ (2)
where $\ell$ denotes $D_{\mathrm{KL}}$ represent the cross-entropy and Kullback-Leibler divergence, respectively. $z_{n}^{(l)}=g(x_{n}^{(l)})$ and $z_{m}^{(u)}=g(x_{m}^{(u)})\in\mathbb{R}^{d}$ are feature representations obtained by the feature extractor $g$. $p_{n}^{(l)}$ and $p_{m}^{(u)}$ are the categorical probability vectors of labeled $x_{n}^{(l)}$ and unlabeled data $x_{m}^{(u)}$, respectively, obtained by teacher VLM encoders $f_{\mathcal{X}}$ and $f_{\mathcal{T}}$. Another well-studied single-head KD baseline is feature distillation, which leverages mean squared error (MSE) loss to directly align feature representations extracted by the student encoder $g$ and the teacher image encoder $f_{\mathcal{X}}$. We defer the details of feature distillation for VLMs to CLIP-KD (Yang et al., 2024a).
2.2 Dual-Head Optimization (DHO)
Figure 3: The average cosine similarity and inner product over 10 datasets.
Gradient conflicts in SHO.
Logit distillation in SHO, as described in Section˜2.1, provides a simple approach for transferring knowledge from VLMs to task-specific models. However, we find that its performance gain is suboptimal, which we attribute to gradient conflicts between the supervised and KD loss signals which hinder effective feature learning (Yu et al., 2020a; Liu et al., 2021; Chen & Er, 2025). As illustrated in and of Fig.˜3, both the classifier weight $W$ of classifier $h$ and the parameter vector $\theta$ of feature extractor $g$ suffer from gradient conflicts: the cosine similarity between their respective gradients turns negative, i.e., $\mathtt{CosSim}(\nabla_{\theta}\mathcal{L}_{\text{CE}},\nabla_{\theta}\mathcal{L}_{\text{KD}})<0$ and $\mathtt{CosSim}(\nabla_{W}\mathcal{L}_{\text{CE}},\nabla_{W}\mathcal{L}_{\text{KD}})<0$, indicating misaligned optimization directions.
To understand this, we first analyze the gradient with respect to (w.r.t) the weight $W$ of a linear head $h$. The gradients w.r.t the classifier weight $W$ are $\nabla_{W}\mathcal{L}_{\text{CE}}=(\hat{p}-y)z^{\top}$ and $\nabla_{W}\mathcal{L}_{\text{KD}}=(\hat{p}-p)z^{\top}$, respectively. Their cosine similarity is proportional to $(\hat{p}-y)^{\top}(\hat{p}-p)\cdot\|z\|^{2}$:
$\mathtt{CosSim}(\nabla_{W}\mathcal{L}_{\text{CE}},\nabla_{W}\mathcal{L}_{\text{KD}})\propto(\hat{p}-y)^{\top}(\hat{p}-p)\cdot\|z\|^{2},$ (3)
which misaligns when $(\hat{p}-y)^{\top}(\hat{p}-p)<0$ due to prediction mismatch, falling below zero during training ( $\times$ in Fig.˜3), leading to gradient conflicts ( in Fig.˜3).
We observe that similar gradient conflicts happen on the parameters $\theta$ in the feature extractor $g$. Let the gradients w.r.t to the feature representation $z$ be $\nabla_{z}\mathcal{L}_{\text{CE}}=W^{\top}(\hat{p}-y)$ and $\nabla_{z}\mathcal{L}_{\text{KD}}=W^{\top}(\hat{p}-p)$, Then, applying the chain rule, the gradients with respect to the feature extractor parameters $\theta$ become $\nabla_{\theta}\mathcal{L}_{\text{CE}}=\nabla_{z}\mathcal{L}_{\text{CE}}\cdot\frac{\partial z}{\partial\theta}$ and $\nabla_{\theta}\mathcal{L}_{\text{KD}}=\nabla_{z}\mathcal{L}_{\text{KD}}\cdot\frac{\partial z}{\partial\theta}$. Similarly, the cosine similarity for the feature extractor parameters becomes:
$\mathtt{CosSim}(\nabla_{\theta}\mathcal{L}_{\text{CE}},\nabla_{\theta}\mathcal{L}_{\text{KD}})\propto(\hat{p}-y)^{\top}W\left(\frac{\partial z}{\partial\theta}\right)^{\top}\frac{\partial z}{\partial\theta}W^{\top}(\hat{p}-p)$ (4)
While the complexity of the Jacobian $\frac{\partial z}{\partial\theta}$ makes it difficult to theoretically guarantee gradient conflicts, we empirically observe that gradient conflicts occur in the parameters $\theta$ of the feature extractor $g$ ( in Fig.˜3), with $\mathtt{CosSim}(\nabla_{\theta}\mathcal{L}_{\text{CE}},\nabla_{\theta}\mathcal{L}_{\text{KD}})<0$ . Thus, we hypothesize that the gradient conflicts in the feature extractor $g$ are propagated from the classification head $h$.
Dual-head architecture.
To mitigate this issue, we propose Dual Head Optimization (DHO) to decouple $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{KD}}$ via two independent prediction heads: $h_{\text{CE}}(z)=W_{\text{CE}}z+b_{\text{CE}}$ and $h_{\text{KD}}(z)=W_{\text{KD}}z+b_{\text{KD}}$, with $W_{\text{CE}},W_{\text{KD}}\in\mathbb{R}^{C\times d}$ and $b_{\text{CE}},b_{\text{KD}}\in\mathbb{R}^{C}$. The corresponding losses are:
$\mathcal{L}_{\text{CE}}$ $=\frac{1}{N}\sum_{n}\ell\left(\sigma(h_{\text{CE}}(z_{n}^{(l)})),y_{n}\right),$ (5)
$\mathcal{L}_{\text{KD}}$ $=\frac{1}{N}\sum_{n}D_{\mathrm{KL}}\left[p_{n}^{(l)}\|\sigma(h_{\text{KD}}(z_{n}^{(l)}))\right]+\frac{1}{M}\sum_{m}D_{\mathrm{KL}}\left[p_{m}^{(u)}\|\sigma(h_{\text{KD}}(z_{m}^{(u)}))\right],$ (6)
where the final loss combines both objectives as $\lambda\mathcal{L}_{\text{CE}}+(1-\lambda)\mathcal{L}_{\text{KD}}$.
Mitigation of gradient conflict in DHO.
In DHO, gradient conflicts in the classification head naturally disappear by decoupling the optimization of $W_{\text{CE}}$ and $W_{\text{KD}}$. This separation enables each head to learn distinct signals without interference, reducing prediction mismatch ( $\times$ of Fig.˜3). Let $\hat{p}_{\text{CE}}{=}\sigma(h_{\text{CE}}(z))$ and $\hat{p}_{\text{KD}}{=}\sigma(h_{\text{KD}}(z))$, then $\nabla_{z}\mathcal{L}_{\text{CE}}{=}W_{\text{CE}}^{\top}(\hat{p}_{\text{CE}}-y)$ and $\nabla_{z}\mathcal{L}_{\text{KD}}{=}W_{\text{KD}}^{\top}(\hat{p}_{\text{KD}}-p)$. The cosine similarity between gradients w.r.t $\theta$ in DHO is defined as:
$\mathtt{CosSim}(\nabla_{\theta}\mathcal{L}_{\text{CE}},\nabla_{\theta}\mathcal{L}_{\text{KD}})\propto(\hat{p}_{\text{CE}}-y)^{\top}W_{\text{CE}}\left(\frac{\partial z}{\partial\theta}\right)^{\top}\frac{\partial z}{\partial\theta}W_{\text{KD}}^{\top}(\hat{p}_{\text{KD}}-p),$ (7)
where we empirically find that gradient conflicts in $\theta$ of the feature extractor $g$ for DHO are also resolved, i.e., maintaining positive gradient alignment throughout training (i.e., Eq.˜7 > 0; as shown in of Fig.˜3). It enables conflict-free representation learning, and leads to better feature representation compared to SHO, which empirically validated by linear evaluation in Section˜3.3. We defer an algorithm that describes the full training procedure of DHO to Alg.˜1.
2.3 Dual-head Interpolation
After training, we find that using only one of the two heads at inference is suboptimal (see Fig.˜6). Motivated by the mixture-of-experts paradigm (Jacobs et al., 1991), we adopt a simple yet effective inference rule for DHO that linearly interpolates the output probability vectors of two heads:
$\hat{p}_{\textbf{{DHO}}}=\alpha\,\sigma\!\big(h_{\text{CE}}(z)\big)+(1-\alpha)\,\sigma\!\big(h_{\text{KD}}(z)/\beta\big),$ (8)
where $\alpha\in[0,1]$ balances the supervised and KD heads, $\beta>0$ is a temperature that softens the KD logits, and the final prediction is $\arg\max_{c\in[C]}\hat{p}_{\textbf{{DHO}}}$. In practice, we tune $\alpha$ and $\beta$ on a validation set to reflect dataset-specific supervision quality and teacher accuracy, allowing the model to weight the more reliable source. See Alg.˜2 for the full inference procedure.
Effect of $\alpha$ and $\beta$.
We now demonstrate the effect of tuning $\alpha$ and $\beta$ using a validation set. Under mild assumptions, DHO $\varepsilon$-approximates SHO in $\ell_{1}$ by setting $\alpha=\lambda$ and $\beta=1$:
Assumption 1 ($\varepsilon$-convergence).
Assume that, after sufficient training, both heads converge to their respective target distributions with $\ell_{1}$-bounded error:
$\sup_{x}\|\sigma(h_{\mathrm{CE}}(z))-y\|_{1}\leq\varepsilon,\quad\sup_{x}\|\sigma\!(h_{\mathrm{KD}}(z))-p\|_{1}\leq\varepsilon,\quad\text{where}\quad\varepsilon\in\mathbb{R}_{>0}.$ (9)
Theorem 1 (Inference equivalence).
Under Assumption˜1, by setting $\alpha=\lambda$ and $\beta=1$, then $\|\hat{p}_{\textbf{{DHO}}}-\hat{p}_{\mathrm{SHO}}\|_{1}\leq\varepsilon,$ where $\hat{p}_{\mathrm{SHO}}$ is the output of SHO optimally trained with $\lambda$.
Details and proofs are deferred to Appendix˜B. Theorem˜1 implies that $\hat{p}_{\text{SHO}}$ trained with any $\lambda$ can be approximated by the dual-head interpolation in Eq.˜8, by setting $\alpha=\lambda$ and $\beta=1$. Here, $\lambda$ is a training hyperparameter of SHO, while $\alpha$ and $\beta$ are inference hyperparameters of DHO, allowing it to emulate SHO hyperparameter tuning without retraining.
Language-aware initialization for VLM students
In the case of VLM-to-VLM distillation, we leverage the text encoder $f_{\mathcal{T}}$ of teachers when initializing the dual heads $h_{\text{CE}}$ and $h_{\text{KD}}$. Following prior work (Li et al., 2022), we initialize the weights as $W_{\text{CE}},W_{\text{KD}}\leftarrow[f_{\mathcal{T}}(t_{1}),\ldots,f_{\mathcal{T}}(t_{C})]^{\top}\in\mathbb{R}^{C\times d}$. We further align the prediction logic of KD head $h_{\text{KD}}$, with the cosine similarity-based approach of the teacher VLMs as follows:
$h_{\text{KD}}=\frac{1}{\zeta}[\mathtt{CosSim}(g(x),w_{1}),\ldots,\mathtt{CosSim}(g(x),w_{C})]^{\top}\in\mathbb{R}^{C},$ (10)
where $w_{c}\in\mathbb{R}^{d}$ denotes the $c$-th row of $W_{\text{KD}}$.
3 Experiments
3.1 Experimental Setups
Datasets.
For in-distribution evaluation, we use ImageNet (Russakovsky et al., 2015) and 10 fine-grained datasets (Fei-Fei et al., 2004; Parkhi et al., 2012; Krause et al., 2013; Nilsback & Zisserman, 2008; Bossard et al., 2014; Maji et al., 2013; Xiao et al., 2010; Cimpoi et al., 2014; Helber et al., 2019; Soomro, 2012). To assess out-of-distribution (OOD) generalization, we use four ImageNet variants (Recht et al., 2019; Wang et al., 2019; Hendrycks et al., 2021a; b). See Appendix˜D for details.
Baselines.
We compare DHO with conventional single-head KD baselines; CE: training only on labeled dataset $\mathcal{D}^{(l)}$ with cross entropy loss (Eq.˜1), KD (logit): on unlabeled dataset $\mathcal{D}^{(u)}$ with logit distillation (Eq.˜2), and KD (feature): on $\mathcal{D}^{(u)}$ with feature distillation (Yang et al., 2024a). We train on both $\mathcal{D}^{(l)}$ and $\mathcal{D}^{(u)}$ with CE+KD (logit) or CE+KD (feature): combining CE with each KD variant using balancing hyperparameter $\lambda$. We also consider dual-head KD approaches SSKD (He et al., 2021) and DHKD (Yang et al., 2024d), though for different purposes as detailed in Appendix˜A.
For in-distribution evaluation on ImageNet with low-shot settings, we compare against self and semi-supervised learning (Chen et al., 2020; Assran et al., 2021; 2022; Cai et al., 2022; Zheng et al., 2023), CLIP-based-training (Li et al., 2022; Liu et al., 2023b), co-training (Rothenberger & Diochnos, 2023), KD (Chen et al., 2020), and zero-shot VLMs (Zhai et al., 2023; Fang et al., 2023).
For OOD evaluation, we compare against VLM adaptation methods, including VPT (Jia et al., 2022), CoOp (Zhou et al., 2022b), PromptSRC (Khattak et al., 2023b), and CasPL (Wu et al., 2025).
Implementation details.
We evaluate across few-shot (1/2/4/8/16-shot) and low-shot (1%/10%) settings, treating remaining data as unlabeled. We adopt Tip-Adapter-F (Zhang et al., 2021) for few-shot teachers, denoting this variant as DHO-F. For a fair comparison, we use the same hyperparameters for DHO and SHO baselines (except $\alpha,\beta$); for other methods, we report the published results. When validation data is unavailable (e.g., ImageNet), we fix $\beta=0.5$ across all settings. We heuristically set $\alpha=0.4$ for zero-shot teachers and $\alpha=0.2$ for few-shot teachers, reflecting the latter’s higher reliability. In low-shot settings, we use $\alpha=0.5$ due to increased label availability. See Appendix˜C for details.
Table 1: Results on ImageNet under few-shot semi-supervision using ResNet-18 and ResNet-50. DHO consistently outperforms all baselines and even the teacher with ResNet-50 (e.g., +0.7/1.9/2.2/3.4/4.4% with a zero-shot teacher; +1.3/1.5/1.4/1.7/1.4% with a few-shot teacher).
ResNet-18 trained from scratch Self-supervised ResNet-50
Method 0-shot 1-shot 2-shot 4-shot 8-shot 16-shot 0-shot 1-shot 2-shot 4-shot 8-shot 16-shot
Single-head KD methods
KD 51.0 - - - - - 61.0 - - - - -
CE - 0.7 1.1 1.8 3.4 8.2 - 11.4 17.3 26.4 36.7 47.0
CE+KD (feature) - 17.1 23.5 28.0 32.2 33.8 - 23.0 32.3 41.3 48.2 54.3
CE+KD (logit) - 50.5 50.6 50.6 51.0 51.2 - 60.4 60.8 61.2 61.6 62.3
Dual-head KD methods
SSKD (He et al., 2021) - 42.5 46.2 48.0 50.6 52.0 - 55.2 58.1 60.0 62.3 64.0
DHKD (Yang et al., 2024d) - 19.7 23.5 23.4 23.7 26.8 - 25.6 34.8 42.7 49.2 55.2
DHO (Ours) - 51.8 52.4 52.6 53.3 54.5 - 61.0 62.1 62.5 63.7 64.7
DHO-F (Ours) - 53.7 54.2 54.8 56.2 57.7 - 62.3 63.1 63.9 65.5 66.8
Teacher Models (Resnet-50)
CLIP 60.3 - - - - - 60.3 - - - - -
Tip-Adapter-F - 61.0 61.6 62.5 63.8 65.4 - 61.0 61.6 62.5 63.8 65.4
Figure 4: Results on 10 datasets under few-shot semi-supervision using ResNet-18 with zero-shot teacher.
Figure 5: Results on 10 datasets using ResNet-18 with either zero- or few-shot teacher.
Table 2: Results on ImageNet and its distribution-shifted variants. $\dagger$: backbone frozen during training. DHO consistently improves VLM adaptation methods and achieves SoTA results on OOD benchmarks.
Method Labeled Data Teacher Model Unlabeled Data Val V2 Sketch R A Avg
CLIP Zero-shot
CLIP - - - 66.7 60.8 46.2 74.0 47.8 59.1
Linear Evaluation
Linear Evaluation 1% - - 72.8 64.2 47.0 74.6 48.6 61.4
DHO 1% - 73.4 65.3 48.3 75.8 49.4 62.4
DHO 1% 74.6 66.4 49.2 76.1 49.8 63.2
Visual Prompt Tuning
VPT 1.25% - - 73.6 64.6 47.7 75.2 48.7 62.0
VPT+DHO 1.25% - 73.6 65.3 48.7 75.9 49.1 62.5
VPT+DHO 1.25% 75.1 66.8 50.0 76.9 50.5 63.9
VLM Text-encoder Prompt Tuning
CoOp 1.25% - - 71.5 64.2 48.0 75.2 49.7 61.7
CoOp+CasPL 1.25% 71.9 64.3 48.3 76.0 - -
CoOp+DHO 1.25% - 72.8 65.5 49.3 76.4 49.5 62.7
CoOp+DHO 1.25% 73.4 66.2 49.5 77.0 50.5 63.3
VLM Multimodal Prompt Tuning
PromptSRC 1.25% - - 71.3 64.4 49.6 77.8 50.9 62.8
PromptSRC+CasPL 1.25% 72.8 65.7 49.7 77.9 - -
PromptSRC+DHO 1.25% - 73.0 65.3 49.5 77.8 51.3 63.4
PromptSRC+DHO 1.25% 73.6 66.1 49.8 78.1 51.0 63.7
3.2 Main Results
Effectiveness of DHO compared to conventional KD baselines.
Tab.˜1 presents results on ImageNet under few-shot semi-supervision using ResNet-18/ResNet-50 student models and ResNet-50 VLM teachers. DHO consistently outperforms all baselines across all settings. Other dual-head methods (i.e., SSKD/DHKD), not designed for few-shot semi-supervision, even underperform CE+KD on logits. Notably, with ResNet-50, DHO outperforms the teacher in every few-shot setting (e.g., +0.7/1.8/2.2/3.4/4.4% or +1.3/1.5/1.4/1.7/1.4% with zero-shot or few-shot teachers).
Next, we evaluate DHO on 10 additional datasets using the ResNet-18 student and ResNet-50 VLM teachers. In Fig.˜4, we observe that DHO also consistently outperforms all baseline methods across 10 datasets, while relative rankings between baselines vary across datasets. Fig.˜5 further demonstrates that the few-shot teacher is more effective than using a zero-shot teacher for DHO. Remarkably, the ResNet-18 student model trained with DHO achieves better performance than the ResNet-50 teacher model in most cases, demonstrating knowledge transfer capability of DHO.
Table 3: Results on ImageNet under low-shot settings. For CT and MCT methods, numbers in parentheses indicate the number of different architectures for co-training.
Method Architecture Params 1% 10%
(M) (%) (%)
Self and Semi-supervised Learning
MSN ViT-B/4 86 75.7 80.2
Semi-ViT ViT-L/14 307 77.3 83.3
Semi-ViT ViT-H/14 632 80.0 84.3
CLIP-based Training
CLIP ViT-B/16 86 74.3 80.4
REACT ViT-B/16 86 76.1 80.8
REACT (Gated-Image) ViT-B/16 129 77.4 81.8
CLIP ViT-L/14 304 80.5 84.7
REACT ViT-L/14 304 81.6 85.1
REACT (Gated-Image) ViT-L/14 380 81.6 85.0
Co-training based Methods
CT Multi-arch (2) 608 80.1 85.1
MCT Multi-arch (2) 608 80.7 85.2
CT Multi-arch (4) 1071 80.0 84.8
MCT Multi-arch (4) 1071 80.5 85.8
Knowledge Distillation
SimCLR v2 distill ResNet-50 (2×+SK) 140 75.9 80.2
SimCLR v2 self-distill ResNet-154 (3×+SK) 795 76.6 80.9
CE + KD (logit) ViT-B/16 86 79.8 80.4
CE + KD (logit) ViT-L/14 304 83.1 83.6
DHO ViT-B/16 86 81.6 82.8
DHO ViT-L/14 304 84.6 85.9
Zero-shot VLMs
SigLIP ViT-SO400M/14 400 83.1 -
DFN ViT-H/14 632 84.4 -
DHO achieves SoTA performance on ImageNet under low-shot settings.
We compare DHO to previous state-of-the-art (SoTA) methods on ImageNet under 1% and 10% labeled data. All results are taken from published papers, except for DHO and CE+KD (logit). As shown in Tab.˜3, DHO with ViT-L/14 surpasses the previous SoTA (e.g., +3.0%/+0.1%), while using fewer parameters (e.g., 76M/767M), on both 1%/10% labeled data. Notably, CE + KD (logit) outperforms semi-supervised methods with the same parameters, demonstrating that they are suboptimal compared to methods leveraging the rich representations of VLMs.
DHO achieves SoTA performance on ImageNet OOD benchmarks.
We evaluate DHO across various VLM adaptation approaches. As shown in Tab.˜2, DHO consistently improves different adaptation methods. Compared to CasPL, DHO exhibits superior performance across both in-distribution (Val) and out-of-distribution (V2/Sketch/R/A) benchmarks, establishing new SoTA OOD results in semi-supervised setting. This suggests that joint training with labeled supervision and teacher distillation provides a more effective strategy than the sequential approach of CasPL, aligning with our hypothesis on DHO for resolving gradient conflicts. See Section˜E.5 for results with fully trained models.
Addtional results.
See Appendix˜E for additional results of MobileNetV2, or comparison to PCGrad (Yu et al., 2020b) and category-aware KD methods (Zhao et al., 2022; Lv et al., 2024).
3.3 Analysis
Table 4: Inference overhead using RTX 4090.
Model Params FLOPs Throughput
(M) (G) (im/s)
ResNet-18 11.69 1.83 3525.7
+ DHO 12.20 (+4.4%) 1.83 (+0.0%) 3518.6 (-0.20%)
ResNet-50 25.56 4.14 1018.4
+ DHO 27.61 (+8.0%) 4.15 (+0.2%) 1016.4 (-0.20%)
Minimal computational overhead of DHO.
As shown in Tab.˜4, DHO introduces negligible computational overhead with minimal parameter increase on the ImageNet with 1000 classes, which can be further reduced down for datasets with fewer classes ($C<1000$). We further provide inference computational overhead for other models in Tab.˜9.
Figure 6: Results of ablation studies on dual-heads interpolation strategy in Eq.˜8 of DHO.
Figure 7: Qualitative results on challenging cases.
Effectiveness of dual-head interpolation.
We evaluate the effectiveness of dual-head interpolation (Eq.˜8) by comparing DHO with CE+KD (logit), CE+KD (feature), and ablations DHO ($h_{\text{CE}}$) and DHO ($h_{\text{KD}}$), which predict using only one head (i.e., $\alpha=1$ or 0). As shown in Fig.˜6, DHO outperforms DHO ($h_{\text{CE}}$) by an average of 1.6% across 11 datasets, with a maximum gain of +3.4% on ImageNet and no degradation on any dataset. Since $\alpha$ and $\beta$ are inference-time hyperparameters, dual-head interpolation introduces minimal overhead while consistently improving or maintaining performance. We also investigate the effectiveness of our adaptive weighting strategy in Section˜E.4. Fig.˜7 illustrates three challenging examples: CE head ($h_{\text{CE}}$) is correct in the first, KD head ($h_{\text{KD}}$) in the second, and both fail in the third case, yet the proposed combined prediction is correct—demonstrating the ability of DHO to resolve individual head failures. See Section˜F.2 for additional analysis on theses challenging examples.
Table 5: Linear evaluation results.
Method Top-1 (%) Top-5 (%)
CE+KD (feature) (Yang et al., 2024a) 62.3 85.0
CE+KD (logit) (Chen et al., 2020) 66.2 88.8
DHO 67.1 89.3
Figure 8: t-SNE visualization.
Enhanced feature representation of DHO.
To validate our claim that mitigating gradient conflicts improves feature representations, we evaluate features using the standard linear evaluation protocol (Chen et al., 2020). We train CE+KD (feature), CE+KD (logit), and DHO under the 16-shot semi-supervised setting on ImageNet, freeze the feature extractor $g$, and train a new prediction linear head $h_{\text{LE}}$ on the top of $g$ using fully labeled data. As shown in Section˜3.3, DHO achieves higher Top-1 and Top-5 accuracy than other methods (e.g., +0.9% and 0.5%, respectively). To further assess feature quality, we visualize embeddings $z$ using t-SNE (Van der Maaten & Hinton, 2008) in Fig.˜8. Compared to the CE+KD (logit) baseline, DHO produces more compact and class-separated feature clusters. These results support our claim that DHO enhances feature representations by mitigating gradient conflicts; this improvement leads to better performance of DHO compared to SHO, as discussed through Tab.˜1, Figs.˜4, 5, 11 and 11.
Table 6: Results on language-aware initialization and KD-head alignment for VLM students on ImageNet with 1% labeled data.
Init ($h_{\text{CE}}$$h_{\text{KD}}$). Align. Accuracy (%)
/ 78.3
/ 78.5 (+0.2)
/ 78.6 (+0.3)
/ 78.7 (+0.4)
Effectiveness of Init. and Align in Section˜2.3.
To assess the effect of the proposed language-aware initialization and KD-head alignment in Section˜2.3, we run ablations that apply language-aware initialization (Init.) to either $h_{\text{CE}}$ or $h_{\text{KD}}$, and optionally enable KD-head alignment (Align.). We evaluate on ImageNet with 1% labels using a ViT-B/16 student and ViT-L/14 teacher for computational efficiency. As shown in Tab.˜6, applying Init. to either head independently improves accuracy (e.g., +0.2/+0.3%), and adding Align. yields additional gains (e.g., +0.4%).
Figure 9: Grid search results for $\alpha$ and $\beta$.
Effect of $\alpha$ and $\beta$.
In Theorem˜1, we show that under mild assumptions (Assumption˜1), DHO $\epsilon$-approximates the SHO baseline, namely CE+KD (logit), in the $\ell_{1}$ norm by setting $\alpha{=}\lambda$ and $\beta{=}1$. This enables DHO to emulate SHO hyperparameter tuning at inference time without retraining. To study how $\alpha$ and $\beta$ affect performance, we visualize a grid search on ImageNet with ResNet-50 under the 16-shot setting. We use ImageNet because it lacks a validation set; thus $\alpha$ and $\beta$ are heuristically set to 0.4 and 0.5 in Section˜3.1. As shown in Fig.˜9, with balanced heads ($\alpha\approx 0.5)$, performance remains stable for $\beta\in[0.1,1]$, and it degrades at extreme $\alpha$ values ($\alpha\approx 0$ or $1$) regardless of $\beta$. The accuracy peaks at 65.5% with balanced heads ($\alpha\approx 0.5$) and a modest temperature ($\beta\approx 0.3$), close to the performance of our heuristic setting (64.7%). Importantly, these gains require no additional training, demonstrating the efficiency of post-training hyperparameter search for the proposed dual-head interpolation in Section˜2.3.
4 Conclusion, Limitation, and Future Work
We identify the fundamental challenge of transferring the zero-/few-shot capabilities of large vision-language models (VLMs) to task-specific models using limited labeled data in semi-supervised settings. Conventional knowledge distillation (KD) methods suffer from gradient conflicts where teacher signals and labeled data signals interfere with each other. We propose DHO (Dual-Head Optimization), a plug-and-play framework using two classification heads with separate objectives. It mitigates gradient conflicts in both heads and feature extractor, improving feature representations, while enabling flexible post-training hyperparameter adjustment via linear combination of outputs. Experiments across 15 datasets show DHO outperforms conventional KD methods, achieving state-of-the-art results on both ImageNet semi-supervised learning with fewer parameters and on out-of-distribution tasks when combined with existing adaptation techniques.
Limitations and future work.
While our main focus is to address the core problem of gradient conflicts arising from general knowledge of VLMs and task-specific patterns from labeled data, we acknowledge several limitations that present opportunities for future research.
We primarily focus on VLMs as general knowledge source, as they provide strong zero-shot and few-shot capabilities in visual recognition tasks. However, we believe the fundamental problem of gradient conflicts between general foundational knowledge and task-specific objectives extends beyond VLMs. This conflict likely emerges in various scenarios where large pre-trained foundation models (such as instruction-tuned language models (Wei et al., 2021; Liu et al., 2023a; Bai et al., 2023; Achiam et al., 2023; Team et al., 2023)) are adapted to specialized downstream tasks. Exploring how DHO performs across diverse foundation models and modalities remains an important direction for future work.
Our implementation is limited to visual recognition tasks, as they represent the most fundamental domain in computer vision and provide an ideal testbed for analyzing gradient conflicts in knowledge transfer. Also, VLMs’ strong zero-shot and few-shot capabilities in visual recognition tasks make them natural candidates for knowledge distillation. However, extending our approach to more complex visual understanding tasks such as object detection and segmentation would be a promising direction with dedicated architectural adaptations.
Reproducibility Statement
We ensure reproducibility by conducting all experiments on publicly available datasets including ImageNet (Russakovsky et al., 2015), Caltech101 (Fei-Fei et al., 2004), and nine other standard benchmarks detailed in Appendix˜D. All experimental configurations are fully specified in Section˜3.1, including exact hyperparameters, learning rate and optimizer settings, model architectures (ResNet-18, ResNet-50, MobileNetV2, ViT-B/16, ViT-L/14), and training procedures with dual-head optimization detailed in Algorithm 1. We use publicly available pre-trained models (CLIP ResNet-50, CLIP ViT variants from OpenAI, DINO ResNet-50) with exact checkpoint specifications provided in Tab.˜7. The inference procedure with hyperparameters $\alpha$ and $\beta$ is fully documented in Algorithm 2, with specific values for each setting ($\alpha=0.4$ for zero-shot, $\alpha=0.2$ for few-shot teachers, $\beta=0.5$ when validation unavailable). We commit to releasing our complete codebase, training scripts, pretrained checkpoints for ImageNet, and evaluation protocols upon acceptance. All experiments are conducted using PyTorch with fixed random seeds on NVIDIA RTX 4090 GPUs (4× for ImageNet, 8× for VLM distillation, single GPU for other benchmarks).
Ethics Statement
Our work presents no new ethical concerns as DHO is a purely technical contribution for knowledge distillation using existing publicly available datasets (ImageNet, Caltech101, and standard computer vision benchmarks) that contain no personally identifiable information. No additional data collection, human subjects research, or sensitive information processing is involved in this work. We acknowledge that vision-language models may contain biases from their pre-training data, which our distillation framework preserves without amplification. The computational requirements vary by dataset scale (single GPU for most benchmarks, 4× GPUs for ImageNet, 8× GPUs for VLM distillation), which remains modest compared to training large vision-language models from scratch, promoting research accessibility while minimizing environmental impact.
Appendix Overview
This appendix provides supplementary material to support the main paper and is organized as follows:
  • Related Work (Appendix˜A) discusses previous work relevant to ours, such as vision-language pre-training, data-limited adaptation of VLMs, knowledge Distillation (KD), and dual head approaches.
  • Theoretical Analysis (Appendix˜B): provides mathematical foundations and theoretical guarantees for our approach.
  • Algorithms and Implementation (Appendix˜C): presents detailed pseudocode (Section˜C.1), implementation specifics (Section˜C.2), and computational overhead analysis (Section˜C.3).
  • Datasets (Appendix˜D): describes the datasets used in our experiments, including statistics and preprocessing details.
  • Additional Experiments (Appendix˜E): presents MobileNet experiments (Section˜E.1), additional results with KD methods (Section˜E.3), additional results with gradient surgery methods (Section˜E.2), additional results with adaptive weighting (Section˜E.4), and results of out-of-distribution generalization with fully-trained models (Section˜E.5).
  • Additional Analyses (Appendix˜F): contains non-linear head design studies (Section˜F.1), and further dual-head investigations (Section˜F.2).
Appendix A Related Work
Vision-language pre-training.
The emergence of vision-language pre-training has marked a significant breakthrough, enabling the use of extensive image-text pairs collected from the web (Wang et al., 2023; Chen et al., 2023) to train powerful vision encoders transferable to various vision tasks (Gan et al., 2022; Zhang et al., 2024b). Early works such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) leveraged contrastive learning techniques to align images and text into a joint representation space, facilitating zero-shot transfer via language prompts. Building on these foundations, subsequent research has focused on improving vision-language models through enhanced training methodologies (Dong et al., 2023; Gao et al., 2022; Yu et al., 2022; Zhai et al., 2023), as well as scaling models and datasets (Yu et al., 2022; Li et al., 2023; Dehghani et al., 2023; Sun et al., 2023a; Cherti et al., 2023; Fang et al., 2023; Sun et al., 2024; Guo et al., 2024) with their zero-shot transfer capabilities (Jia et al., 2021; Zhai et al., 2022; Pham et al., 2023; Liu et al., 2023b). In contrast, our work focuses specifically on target tasks with compact models, aiming to distill knowledge from these large VLMs effectively.
Data-limited adaptation of VLMs.
To preserve pretrained semantic features of VLMs during adaptation with limited data, several approaches have been proposed. Prompt tuning (Lester et al., 2021), initially designed for language models, has been successfully extended to vision tasks. Various methods (Jia et al., 2022; Zhou et al., 2022b; a; Khattak et al., 2023a; Zhu et al., 2023; Khattak et al., 2023b; Menghini et al., 2023; Zhao et al., 2024; Roy & Etemad, 2023; Zhang et al., 2024a; Lafon et al., 2025) have demonstrated the effectiveness of training learnable prompts while keeping the base model frozen. Adapters (Gao et al., 2024; Zhang et al., 2021; Yu et al., 2023b; Silva-Rodriguez et al., 2024) provide an alternative approach by introducing lightweight, trainable modules while maintaining the pre-trained backbone intact. LP++ (Huang et al., 2024) has shown that simple linear layers can effectively adapt CLIP representations in data-limited settings. Note that our work is orthogonal to these approaches: we aim to distill the knowledge of pretrained VLMs into compact models under data-scarce scenarios, making these adaptation methods complementary and applicable to both teacher VLMs in our framework and student models when they are also VLMs.
Knowledge Distillation (KD; Hinton, 2015) enables transferring knowledge from large teacher models to compact student architectures, particularly in data-constrained settings. Researchers have explored synthetic data generation (Lopes et al., 2017; Kimura et al., 2018; Nayak et al., 2019; Yoo et al., 2019; Chen et al., 2019; Yin et al., 2020; Fang et al., 2021a; Nguyen et al., 2022; Patel et al., 2023; Yu et al., 2023a; Liu et al., 2024; Tran et al., 2024; Wang et al., 2024), semi-supervised (Chen et al., 2020; He et al., 2021; Du et al., 2023; Yang et al., 2024b), and unsupervised KD using self-supervised teachers (Fang et al., 2021c; Abbasi Koohpayegani et al., 2020; Navaneet et al., 2021; Wang et al., 2022; Xu et al., 2021; Singh & Wang, 2025). In the VLM domain, recent works (Fang et al., 2021b; Wu et al., 2023; Sun et al., 2023b; Yang et al., 2024a; Vasu et al., 2024; Udandarao et al., 2024; Yang et al., 2024c) distill from large-scale vision-language models into smaller architectures, often using transductive (Kim et al., 2024; Chen et al., 2024) or multi-stage unsupervised strategies (Vemulapalli et al., 2024; Wu et al., 2025; Mistretta et al., 2025). Meanwhile, KD remains challenging due to numerous issues, including model capacity gaps (Cho & Hariharan, 2019; Mirzadeh et al., 2020; Zhu & Wang, 2021; Huang et al., 2022; Li et al., 2024) and inconsistencies between soft and hard targets (Zhang et al., 2023). These challenges are further complicated by misalignment between labeled data and foundational knowledge, especially in few-shot learning scenarios where limited labeled examples may not fully capture the rich semantic understanding of foundation models.
Dual head approaches.
We also consider existing dual-head KD approaches: SSKD (He et al., 2021) uses dual-heads, with each head trained on labeled and unlabeled sets respectively, assuming different data distributions between them, and DHKD (Yang et al., 2024d) introduces binary KD loss working on logits before softmax to prevent neural collapse (Papyan et al., 2020), but both methods inference using only single prediction head $h_{\text{CE}}$. While these previous KD methods adopt dual-head architectures, they do not target distillation from foundation models or combine predictions at inference. In contrast, DHO addresses gradient conflicts in this setting and leverages dual-head aggregation at inference to enhance performance with minimal hyperparameter tuning costs.
Appendix B Theoretical Analysis
In this section, we provide a theoretical analysis of our Dual-Head Optimization (DHO) framework. We establish that DHO effectively addresses single-head logit distillation Hinton (2015); Chen et al. (2020) by decoupling conflicting gradients through specialized heads during training. We prove that post-training, the optimal prediction from our dual-head model—formulated as a weighted combination of the heads’ outputs—is mathematically equivalent to the optimal solution of conventional single-head distillation. This equivalence provides theoretical justification for our approach while eliminating gradient conflicts. Furthermore, DHO enables efficient adaptation to various datasets through tunable hyperparameters ($\alpha$ and $\beta$) without requiring model retraining. Note that in this section we slightly abuse the notation of the main paper for clarity, e.g., we denote $p_{\tau}$ as teacher predictions with temperature scaling $\tau$.
B.1 Single-Head Optimization
We begin by considering two target probability distributions: the ground truth label distribution $y$ and the teacher’s softened distribution $p_{\tau}$ for input $x\in\mathcal{X}$, where:
  • $y$ represents the ground truth label distribution, typically one-hot encoded vectors where $y_{c}=1$ for the true class $c$ and 0 elsewhere
  • $p_{\tau}$ denotes the teacher’s softened distribution with temperature scaling: $p_{\tau}=\sigma(z_{t}/\tau)$, where $z_{t}$ represents the teacher’s logits and $\sigma$ is the softmax function
Theorem 2 (Optimal Distribution for Single-Head Optimization).
The distribution $\hat{p}^{*}$ that minimizes the weighted combination of cross-entropy loss with respect to $y$ and Kullback-Leibler divergence with respect to $p_{\tau}$:
[…]
is given by the weighted arithmetic mean:
$\hat{p}^{*}=\lambda y+(1-\lambda)p_{\tau}$ (12)
where $\lambda\in[0,1]$ is the weighting hyperparameter.
Proof.
We begin by expanding the objective function:
[…]
$=-\lambda\sum_{c=1}^{C}y_{c}\log\hat{p}_{c}+(1-\lambda)\sum_{c=1}^{C}p_{\tau,c}\log\frac{p_{\tau,c}}{\hat{p}_{c}}$ (14)
$=-\lambda\sum_{c=1}^{C}y_{c}\log\hat{p}_{c}+(1-\lambda)\sum_{c=1}^{C}p_{\tau,c}\log p_{\tau,c}-(1-\lambda)\sum_{c=1}^{C}p_{\tau,c}\log\hat{p}_{c}$ (15)
$=-\sum_{c=1}^{C}[\lambda y_{c}+(1-\lambda)p_{\tau,c}]\log\hat{p}_{c}+(1-\lambda)\sum_{c=1}^{C}p_{\tau,c}\log p_{\tau,c}$ (16)
Since the last term is constant with respect to $\hat{p}$, the optimization problem reduces to minimizing:
$\mathcal{L}^{\prime}(\hat{p})=-\sum_{c=1}^{C}[\lambda y_{c}+(1-\lambda)p_{\tau,c}]\log\hat{p}_{c}$ (17)
Subject to the probability constraints:
$\sum_{c=1}^{C}\hat{p}_{c}=1,\quad\hat{p}_{c}\geq 0\quad\forall c\in\{1,2,\ldots,C\}$ (18)
Applying the method of Lagrange multipliers with multiplier $\mu$:
$\mathcal{L}(\hat{p},\mu)=-\sum_{c=1}^{C}[\lambda y_{c}+(1-\lambda)p_{\tau,c}]\log\hat{p}_{c}+\mu\left(\sum_{c=1}^{C}\hat{p}_{c}-1\right)$ (19)
Taking the partial derivative with respect to $\hat{p}_{c}$ and setting it to zero:
$\frac{\lambda y_{c}+(1-\lambda)p_{\tau,c}}{\hat{p}_{c}}+\mu=0$ (20)
Solving for $\hat{p}_{c}$:
$\hat{p}_{c}=\frac{\lambda y_{c}+(1-\lambda)p_{\tau,c}}{\mu}$ (21)
Using the constraint $\sum_{c=1}^{C}\hat{p}_{c}=1$, and observing that $\sum_{c=1}^{C}y_{c}=1$ and $\sum_{c=1}^{C}p_{\tau,c}=1$ (both being probability distributions):
$\sum_{c=1}^{C}\hat{p}_{c}=\sum_{c=1}^{C}\frac{\lambda y_{c}+(1-\lambda)p_{\tau,c}}{\mu}=1$ (22)
$\frac{1}{\mu}\sum_{c=1}^{C}[\lambda y_{c}+(1-\lambda)p_{\tau,c}]=1$ (23)
$\frac{1}{\mu}[\lambda\sum_{c=1}^{C}y_{c}+(1-\lambda)\sum_{c=1}^{C}p_{\tau,c}]=1$ (24)
$\frac{1}{\mu}[\lambda+(1-\lambda)]=1$ (25)
$\mu=1$ (26)
Therefore, the optimal solution is:
$\hat{p}^{*}_{c}=\lambda y_{c}+(1-\lambda)p_{\tau,c}$ (27)
This weighted arithmetic mean of the two target distributions is the optimal solution that minimizes our objective function. ∎
B.2 Dual-Head Optimization
In our proposed Dual-Head Optimization (DHO) framework, we extract shared features $g(x)$ from input $x$ and apply two specialized classification heads:
  • $h_{\text{CE}}(z)=W_{\text{CE}}z+b_{\text{CE}}$: optimized exclusively to match ground truth labels using cross-entropy loss $\ell(\sigma(h_{\text{CE}}(z)),y)$
  • $h_{\text{KD}}(z)=W_{\text{KD}}z+b_{\text{KD}}$: optimized exclusively to match teacher predictions using KL divergence $D_{\mathrm{KL}}(p_{\tau}\|\sigma(h_{\text{KD}}(z)/\beta))$
where $z=g(x)$ is the feature representation, and the parameter $\beta$ controls the temperature during inference, while a fixed temperature of 1 is used during training of the knowledge distillation head.
Assumption 2 ($\varepsilon$-Convergence).
We assume that after sufficient training, both heads have converged to their respective target distributions with bounded error:
$\sup_{x}\|\sigma(h_{\text{CE}}(z))-y\|_{1}\leq\varepsilon,\quad\sup_{x}\|\sigma(h_{\text{KD}}(z)/\beta)-p_{\tau}\|_{1}\leq\varepsilon$ (28)
where $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm and $\varepsilon>0$ is a small constant.
Theorem 3 (Inference Equivalence Under $\varepsilon$-Convergence).
Under Assumption˜2, by combining the outputs of both heads as:
$\hat{p}_{\textbf{{DHO}}}=\alpha\cdot\sigma(h_{\text{CE}}(z))+(1-\alpha)\cdot\sigma(h_{\text{KD}}(z)/\beta),\quad\text{where }\alpha=\lambda$ (29)
we obtain a prediction that approximates the optimal single-head solution with bounded error:
$\|\hat{p}_{\textbf{{DHO}}}-\hat{p}^{*}\|_{1}\leq\varepsilon$ (30)
Proof.
We analyze the $\ell_{1}$ distance between the DHO prediction and the optimal solution:
$\|\hat{p}_{\textbf{{DHO}}}-\hat{p}^{*}\|_{1}$ $=\|\alpha\cdot\sigma(h_{\text{CE}}(z))+(1-\alpha)\cdot\sigma(h_{\text{KD}}(z)/\beta)-\lambda y-(1-\lambda)p_{\tau}\|_{1}$ (31)
$=\|\lambda(\sigma(h_{\text{CE}}(z))-y)+(1-\lambda)(\sigma(h_{\text{KD}}(z)/\beta)-p_{\tau})\|_{1}$ (32)
$\leq\lambda\|\sigma(h_{\text{CE}}(z))-y\|_{1}+(1-\lambda)\|\sigma(h_{\text{KD}}(z)/\beta)-p_{\tau}\|_{1}$ (33)
$\leq\lambda\varepsilon+(1-\lambda)\varepsilon=\varepsilon$ (34)
where we applied the triangle inequality for the $\ell_{1}$ norm and used Assumption˜2.
Therefore, we have established that:
$\hat{p}_{\textbf{{DHO}}}\approx_{\varepsilon}\hat{p}^{*}$ (35)
where $\approx_{\varepsilon}$ denotes approximation with $\ell_{1}$ error bound $\varepsilon$. ∎
Lemma 1 (Temperature Matching via KL Divergence).
Assume the knowledge distillation head is trained to minimize KL divergence with respect to the teacher’s predictions at temperature 1, such that:
$D_{\mathrm{KL}}(p_{1}\|\sigma(h_{\text{KD}}(z)))\leq\delta$ (36)
Then, setting the temperature parameter $\beta=\tau$ at inference time allows the KD head to approximate the teacher’s prediction at temperature $\tau$ with error bound:
$\|\sigma(h_{\text{KD}}(z)/\beta)-p_{\tau}\|_{1}\leq\sqrt{2\delta}$ (37)
Proof.
When logits are properly scaled and under appropriate conditions of the softmax function, we can reasonably approximate:
$D_{\mathrm{KL}}(p_{\tau}\|\sigma(h_{\text{KD}}(z)/\tau))\approx D_{\mathrm{KL}}(p_{1}\|\sigma(h_{\text{KD}}(z)))\leq\delta$ (38)
Applying Pinsker’s inequality, which establishes a relationship between KL divergence and the L1 norm difference between probability distributions:
$\|p_{\tau}-\sigma(h_{\text{KD}}(z)/\tau)\|_{1}\leq\sqrt{2D_{\mathrm{KL}}(p_{\tau}\|\sigma(h_{\text{KD}}(z)/\tau))}\leq\sqrt{2\delta}$ (39)
To ensure $\varepsilon$-convergence between the KD head at temperature $\tau$ and the teacher’s prediction at temperature $\tau$, it is sufficient to guarantee:
$\sqrt{2\delta}\leq\varepsilon\Rightarrow\delta\leq\frac{\varepsilon^{2}}{2}$ (40)
Corollary 1 (Optimal DHO Configuration).
With proper training ensuring $\varepsilon$-convergence of both heads, dual-head optimization with temperature parameter $\beta=\tau$ and mixing parameter $\alpha=\lambda$ approximates the optimal single-head objective with error bounded by $\varepsilon$:
$\hat{p}_{\textbf{{DHO}}}\approx_{\varepsilon}\hat{p}^{*}=\lambda y+(1-\lambda)p_{\tau}$ (41)
This demonstrates that our DHO approach achieves the same theoretical optimality as SHO.
Appendix C Algorithms and Implementation
C.1 Pseudocode
We present the pseudocode for DHO in Algs.˜1 and 2 for training and inference, respectively.
Algorithm 1 DHO Training with zero-shot CLIP (Radford et al., 2021) teacher
1:Input: labeled set $\mathcal{D}^{(l)}=\{(x^{(l)}_{i},y_{i})\}_{i=1}^{N}$, unlabeled set $\mathcal{D}^{(u)}=\{x^{(u)}_{j}\}_{j=1}^{M}$,
2:Input: student feature extractor $g$, prediction heads $h_{\text{CE}},h_{\text{KD}}$, teacher encoders $f_{\mathcal{X}},f_{\mathcal{T}}$,
3:Input: prompt template “A photo of [CLASS]”, temperature scaling factors $\zeta,\eta$,
4:Input: balancing hyperparameter $\lambda$,
5:Input: supervised mini-batch size $B$, and unsupervised mini-batch size $B^{\prime}$.
6: while not converged do
7:  Sample mini-batch $\mathcal{B}^{(l)}=\{(x_{b}^{(l)},y_{b})\}_{b=1}^{B}$ from $\mathcal{D}^{(l)}$, $\mathcal{B}^{(u)}=\{x_{b^{\prime}}^{(u)}\}_{b^{\prime}=1}^{B^{\prime}}$ from $\mathcal{D}^{(l)}\cup\mathcal{D}^{(u)}$.
8:  // Process labeled data
9:  for each $(x_{b}^{(l)},y_{b})\in\mathcal{B}^{(l)}$ do
10:   $z_{b}^{(l)}\leftarrow g(x_{b}^{(l)})$
11:   $\hat{p}_{\text{CE},b}^{(l)}\leftarrow\sigma(h_{\text{CE}}(z_{b}^{(l)}))$
12:   $\hat{p}_{\text{KD},b}^{(l)}\leftarrow\sigma(\frac{1}{\eta}h_{\text{KD}}(z_{b}^{(l)}))$
13:   $p_{b}^{(l)}\leftarrow\sigma\left(\frac{1}{\zeta\cdot\eta}[\mathtt{CosSim}(f_{\mathcal{X}}(x_{b}^{(l)}),f_{\mathcal{T}}(t_{1})),\ldots,\mathtt{CosSim}(f_{\mathcal{X}}(x_{b}^{(l)}),f_{\mathcal{T}}(t_{C}))]^{\top}\right)$
14:  end for
15:  // Process unlabeled data
16:  for each $x_{b^{\prime}}^{(u)}\in\mathcal{B}^{(u)}$ do
17:   $z_{b^{\prime}}^{(u)}\leftarrow g(x_{b^{\prime}}^{(u)})$
18:   $\hat{p}_{\text{KD},b^{\prime}}^{(u)}\leftarrow\sigma(\frac{1}{\eta}h_{\text{KD}}(z_{b^{\prime}}^{(u)}))$
19:   $p_{b^{\prime}}^{(u)}\leftarrow\sigma\left(\frac{1}{\zeta\cdot\eta}[\mathtt{CosSim}(f_{\mathcal{X}}(x_{b^{\prime}}^{(u)}),f_{\mathcal{T}}(t_{1})),\ldots,\mathtt{CosSim}(f_{\mathcal{X}}(x_{b^{\prime}}^{(u)}),f_{\mathcal{T}}(t_{C}))]^{\top}\right)$
20:  end for
21:  // Compute losses and update
22:  $\mathcal{L}_{\text{CE}}\leftarrow\frac{1}{B}\sum_{b=1}^{B}\ell(\hat{p}_{\text{CE},b}^{(l)},y_{b})$
23:  $\mathcal{L}_{\text{KD}}\leftarrow\frac{1}{B}\sum_{b=1}^{B}D_{\mathrm{KL}}(\hat{p}_{\text{KD},b}^{(l)}||p_{b}^{(l)})+\frac{1}{B^{\prime}}\sum_{b^{\prime}=1}^{B^{\prime}}D_{\mathrm{KL}}(\hat{p}_{\text{KD},b^{\prime}}^{(u)}||p_{b^{\prime}}^{(u)})$
24:  $\mathcal{L}\leftarrow\lambda\mathcal{L}_{\text{CE}}+(1-\lambda)\mathcal{L}_{\text{KD}}$
25:  Update parameters of $g$, $h_{\text{CE}}$, $h_{\text{KD}}$ using $\nabla\mathcal{L}$
26: end while
Algorithm 2 Dual-Head Optimization Inference
1:Input: an image $x$, feature extractor $g$, prediction heads $h_{\text{CE}},h_{\text{KD}}$, linear coefficient $\alpha$, temperature scaling $\beta$
[…]
7:Return: $\hat{y}$
C.2 Implementation Details
Tab.˜7 provides a comprehensive overview of the implementation details for our experiments on 1) few-shot semi-supervised settings on ImageNet and 10 datasets, 2) low-shot semi-supervised settings on ImageNet, and 3) VLM-based adaptation methods.
Table 7: Implementation details for our experiments across different settings.
Few-shot Semi-supervised Settings on ImageNet
Model Configuration Student Training Details
Student: ResNet18 (He et al., 2016) from scratch or ResNet50 from DINO (Caron et al., 2021) Input size: $\times224$ Zero-shot Teacher: ResNet50 from CLIP (Radford et al., 2021) Few-shot Teacher: ResNet50 from Tip-Adapter-F (Zhang et al., 2021) Teacher input size: $\times224$ labeled data: $K\in\{1,2,4,8,16\}$ shots $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: $\alpha=0.4$, $\beta=0.5$ (zero-shot); $\alpha=0.2$, $\beta=0.5$ (few-shot) Epochs: 20 Optimizer: AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) Learning rate: $1\times 10^{-3}$, weight decay: $1\times 10^{-2}$ Batch size: 512 (labeled: 256, unlabeled: 256) Scheduler: Cosine decay without warmup Augmentation: Random crops (x0.5-1.0), horizontal flips
Few-shot Semi-supervised Settings on 10 Fine-Grained Datasets
Model Configuration Student Training Details
Student: ResNet18 (He et al., 2016) or MobileNet (Sandler et al., 2018) pre-trained on ImageNet under supervision Input size: $\times224$ Zero-shot Teacher: ResNet50 from CLIP (Radford et al., 2021) Few-shot Teacher: ResNet50 from Tip-Adapter-F (Zhang et al., 2021) Teacher input size: $\times224$ labeled data: $K\in\{1,2,4,8,16\}$ shots $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: determined by validation Epochs: 200 Optimizer: AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) Learning rate: $1\times 10^{-3}$, weight decay: $1\times 10^{-2}$ Batch size: 128 (labeled: 64, unlabeled: 64) Scheduler: Cosine decay without warmup Augmentation: Random crops (x0.5-1.0), horizontal flips
Low-shot Semi-supervised Settings on ImageNet
Model Configuration Student Training Details
Student: CLIP ViT-B/16 or ViT-L/14 (Radford et al., 2021) Input size: $\times224$ (ViT-B/16) or $\times336$ (ViT-L/14) Zero-shot Teacher: CLIP ViT-L/14 or ViT-H/14 (Fang et al., 2023) Teacher input size: $\times336$ (ViT-L/14) or $\times378$ (ViT-H/14) Few-shot Teacher: N/A labeled data: 1% ($\frac{N}{N+M}\approx 0.01$) or 10% ($\frac{N}{N+M}\approx 0.1$) of training data $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: $\alpha=0.5$, $\beta=0.5$ Epochs: 32 Optimizer: AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) Learning rate: $5\times 10^{-5}$, weight decay: $5\times 10^{-2}$ Batch size: 512 (labeled: 256, unlabeled: 256) Scheduler: Cosine warmup decay (5000 steps) Augmentation: Random crops (x0.5-1.0), horizontal flips
Tab.˜8 provides the implementation details for our out-of-distribution generalization experiments on 1) full training model evaluation and 2) adaptation methods including linear evaluation, visual prompt tuning, and VLM-based methods.
Table 8: Implementation details for our out-of-distribution generalization experiments.
Full Training
Model Configuration Training Details
Student: CLIP ViT-B/16 (Radford et al., 2021) Student input size: $\times224$ Zero-shot Teacher: CLIP ViT-L/14 (Radford et al., 2021) Teacher input size: $\times336$ Labeled data: 1% and 10% ImageNet $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: $0.5$ and 11 Epochs: 32 Optimizer: AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) Learning rate: $5\times 10^{-5}$, weight decay: $5\times 10^{-2}$ Batch size: 512 (labeled: 256, unlabeled: 256) Scheduler: Cosine warmup decay (5000 steps) Augmentation: Random crops (x0.5-1.0), horizontal flips
Adaptation Methods (Linear Evaluation & Visual Prompt Tuning)
Method Configuration Training Details
Linear evaluation (Caron et al., 2021) Visual prompt tuning (Jia et al., 2022) Frozen backbone: CLIP ViT-B/16 (Radford et al., 2021) Input size: $\times224$ Zero-shot Teacher: CLIP ViT-L/14 (Radford et al., 2021) Teacher input size: $\times336$ Labeled data: 1% and 10% ImageNet $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: $0.5$ and 11 Epochs: 20 Optimizer: AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) Learning rate: $5\times 10^{-5}$, weight decay: $5\times 10^{-2}$ Batch size: 512 (labeled: 256, unlabeled: 256) Scheduler: Cosine warmup decay (5000 steps) Augmentation: Random crops (x0.5-1.0), horizontal flips
Adaptation Methods (Prompt Tuning)
Method Configuration Training Details
Prompt tuning: CoOp (Zhou et al., 2022b), PromptSRC (Khattak et al., 2023b) Frozen backbone: CLIP ViT-B/16 (Radford et al., 2021) Input size: $\times224$ Zero-shot Teacher: CLIP ViT-L/14 (Radford et al., 2021) Teacher input size: $\times336$ Labeled data: 1% and 10% ImageNet $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: $0.5$ and 11 Prompt tuning: Following PromptSRC (Khattak et al., 2023b) configurations Comparison: CasPL (Wu et al., 2025) with domain-specific unlabeled data
C.3 Computational Costs
Inference overhead of DHO over SHO.
Tab.˜9 presents computational overheads at inference time introduced by DHO over SHO for all the architectures in this paper, such as MobileNetV2 (Sandler et al., 2018), ResNet-18 (He et al., 2016), ResNet-50 (He et al., 2016), ViT-B/16 (Dosovitskiy, 2020), and ViT-L/16 (Dosovitskiy, 2020).
Table 9: Inference overhead using RTX 4090 across different architecture.
Model Params (M) FLOPs (G) Throughput (im/s)
MobileNetV2 3.50 0.33 2978.4
+ DHO 4.79 (+36.5%) 0.34 (+3.0%) 2971.2 (-0.24%)
ResNet-18 11.69 1.83 3525.7
+ DHO 12.20 (+4.4%) 1.83 (+0.0%) 3518.6 (-0.20%)
ResNet-50 25.56 4.14 1018.4
+ DHO 27.61 (+8.0%) 4.15 (+0.2%) 1016.4 (-0.19%)
ViT-B/16 86.57 16.87 290.2
+ DHO 87.34 (+0.9%) 16.87 (+0.0%) 290.1 (-0.02%)
ViT-L/16 304.33 59.70 255.1
+ DHO 305.35 (+0.3%) 59.70 (+0.0%) 255.6 (+0.18%)
Table 10: Training time and hardware.
Student Teacher Training Time Hardware
ResNet-18 ResNet-50 $\approx$ 6 hours 4× RTX 4090
ResNet-50 ResNet-50 $\approx$ 8 hours 4× RTX 4090
ViT-B/16 ViT-L/14 $\approx$ 28 hours 8× RTX 4090
ViT-B/16 ViT-H/14 $\approx$ 40 hours 8× RTX 4090
ViT-L/14 ViT-H/14 $\approx$ 80 hours 8× RTX 4090
Training time and hardware requirements.
Tab.˜10 presents the training time required for our experiments. For VLM distillation experiments, which represent the most resource-intensive component of our work, we used 8× NVIDIA RTX 4090 GPUs. The ViT-H/14 to ViT-L/14 distillation required approximately 80 hours, while the ViT-H/14 to ViT-B/16 and ViT-L/14 to ViT-B/16 distillations required approximately 40 and 28 hours, respectively. For the ViT-H/14 to ViT-L/14 distillation, we implemented gradient accumulation with 4 steps and mixed precision training (Micikevicius et al., 2017) to optimize computational efficiency. For ImageNet experiments, we used 4× NVIDIA RTX 4090 GPUs, with ResNet-18 and ResNet-50 models requiring approximately 6 and 8 hours of training time, respectively. We provide these details to facilitate reproduction of our results and to give researchers a clear understanding of the computational resources needed to implement our approach at scale.
Inference overhead improvements with ToMe.
To further improve the computational efficiency of our approach, we explored integrating Token Merging (ToMe) (Bolya et al., 2022) with DHO. ToMe is a technique that reduces the number of tokens in ViTs by merging similar tokens to improve the efficiency of ViTs. Tab.˜11 shows that combining DHO with ToMe significantly reduces computational costs with minimal impact on performance.
Table 11: Performance and inference overhead of DHO with Token Merging (ToMe) on ImageNet under low-shot semi-supervised settings using RTX 4090.
Method Labeled Accuracy (%) Params (M) FLOPs (G) Throughput (im/s)
DHO 1% 81.6 87.22 17.58 243.35
DHO + ToMe 1% 81.4 (–0.2) 87.22 13.12 (–25.4%) 323.39 (+32.9%)
DHO 10% 82.8 87.22 17.58 238.11
DHO + ToMe 10% 82.5 (–0.3) 87.22 13.12 (–25.4%) 308.49 (+29.6%)
Appendix D Datasets
Table 12: Overview of datasets used in our experiments, organized into three categories: (top) standard classification datasets, (middle) ImageNet, and (bottom) ImageNet variants for out-of-distribution (OOD) evaluation. For few-shot semi-supervised learning experiments, we report both the absolute number of labeled samples and their percentage relative to the full training set.
Dataset # Classes # Train # Val # Test # Labeled (1-shot) # Labeled (16-shot)
Fine-grained 10 Datasets
Caltech101 (Fei-Fei et al., 2004) 100 4,128 1,649 2,465 100 (2.42%) 1,600 (38.76%)
OxfordPets (Parkhi et al., 2012) 37 2,944 736 3,669 37 (1.26%) 592 (20.11%)
StanfordCars (Krause et al., 2013) 196 6,509 1,635 8,041 196 (3.01%) 3,136 (48.18%)
Flowers102 (Nilsback & Zisserman, 2008) 102 4,093 1,633 2,463 102 (2.49%) 1,632 (39.87%)
Food101 (Bossard et al., 2014) 101 50,500 20,200 30,300 101 (0.20%) 1,616 (3.20%)
FGVCAircraft (Maji et al., 2013) 100 3,334 3,333 3,333 100 (3.00%) 1,600 (48.00%)
SUN397 (Xiao et al., 2010) 397 15,880 3,970 19,850 397 (2.50%) 6,352 (40.00%)
DTD (Cimpoi et al., 2014) 47 2,820 1,128 1,692 47 (1.67%) 752 (26.67%)
EuroSAT (Helber et al., 2019) 10 13,500 5,400 8,100 10 (0.07%) 160 (1.19%)
UCF101 (Soomro, 2012) 101 7,639 1,898 3,783 101 (1.32%) 1,616 (21.15%)
Coarse-grained Dataset
ImageNet (Russakovsky et al., 2015) 1,000 1.28M - 50,000 1,000 (0.08%) 16,000 (1.25%)
ImageNet OOD Variants
ImageNet-V2 (Recht et al., 2019) 1,000 - - 10,000 - -
ImageNet-Sketch (Wang et al., 2019) 1,000 - - 50,889 - -
ImageNet-A (Hendrycks et al., 2021b) 200 - - 7,500 - -
ImageNet-R (Hendrycks et al., 2021a) 200 - - 30,000 - -
We evaluated our approach on 11 diverse datasets, with ImageNet (Russakovsky et al., 2015) serving as our primary benchmark. The datasets span general object recognition (Russakovsky et al., 2015; Fei-Fei et al., 2004), fine-grained classification tasks (vehicles (Krause et al., 2013; Maji et al., 2013), natural entities (Nilsback & Zisserman, 2008; Parkhi et al., 2012; Bossard et al., 2014)), and specialized domains (scenes (Xiao et al., 2010), textures (Cimpoi et al., 2014), remote sensing (Helber et al., 2019), and human actions (Soomro, 2012)). Additionally, we conduct experiments on four out-of-distribution test sets to further validate the model’s generalization capabilities. To assess our model’s robustness to distribution shifts, we evaluate it on several challenging variants of ImageNet: ImageNet-v2 (Recht et al., 2019), ImageNet-Sketch (Wang et al., 2019), ImageNet-A (Hendrycks et al., 2021b), and ImageNet-R (Hendrycks et al., 2021a).
We summarize the overview of datasets used in Tab.˜12, these datasets exhibit diversity in their characteristics, with varying numbers of classes and samples per dataset. This diversity enabled us to thoroughly validate our method across different few-shot semi-supervised learning scenarios by systematically varying the ratios between labeled and unlabeled samples.
Appendix E Additional Experiments
E.1 Experiments on MobileNet
Figure 10: Results on 10 datasets under few-shot semi-supervision using MobileNetV2 with zero-shot teacher (Radford et al., 2021).
Figure 11: Results on 10 datasets using MobileNetV2 with either zero- or few-shot teacher (Zhang et al., 2021).
To demonstrate the versatility of our DHO approach beyond ResNet (He et al., 2016) and ViT (Dosovitskiy, 2020) architectures, we extended our experiments to the MobileNetV2 (Sandler et al., 2018), which is specifically designed for real-world applications with compact models. We maintained identical experimental settings as described in Section˜C.2, using MobileNetV2 as the student model while distilling from CLIP ResNet50.
As illustrated in Fig.˜11, our DHO consistently outperforms all single-head baseline methods, demonstrating its effectiveness on lightweight model architectures along with ResNet18. Furthermore, Fig.˜11 reveals patterns similar to our ResNet18 experiments regarding few-shot integration. Our method successfully incorporates few-shot teacher knowledge, although we observe that the few-shot teacher does not consistently yield improvements over the zero-shot teacher. Notably, our distilled MobileNetV2 model sometimes achieves superior performance to the zero- and few-shot teachers (ResNet-50) despite having significantly fewer parameters. This pattern of outperforming both zero-shot and few-shot teachers mirrors the observations from our main experiments, further validating the effectiveness of our approach across different architectural families.
E.2 Additional Results with Gradient Surgery Methods
In this section, we present experimental results comparing DHO with PCGrad (Yu et al., 2020b), a well-known gradient surgery method from the multi-task learning literature. PCGrad addresses gradient conflicts by projecting conflicting gradients to resolve conflicts post hoc.
We compare DHO with PCGrad using 8 of the 10 datasets (without Food101, Sun397) with ResNet-18 as the student model. The results are shown in Tab.˜13.
Table 13: Performance comparison of PCGrad with DHO on 8 datasets (average) using ResNet-18.
Method 1-shot 2-shot 4-shot 8-shot 16-shot
CE+KD (logit) 56.2 58.9 61.7 65.8 69.8
PCGrad 56.8 60.3 62.1 67.5 71.6
DHO 60.1 63.7 70.2 75.4 79.1
We observe that PCGrad improves over CE+KD (logit), but still underperforms compared to DHO. Moreover, PCGrad incurs additional memory and computational costs ($\mathcal{O}(|\theta|)$) due to gradient projection and storage, whereas DHO remains lightweight ($\mathcal{O}(d\times C)$) and simple to implement.
The superior performance of DHO can be attributed to its approach of avoiding gradient conflicts at the source by isolating the learning dynamics of each objective via dual heads, rather than resolving conflicts after they arise. Additionally, DHO provides dynamic interpolation capability at inference time between supervised and distillation signals, which standard gradient-based methods do not offer.
E.3 Additional Results with KD Methods
In this section, we present experimental results combining our DHO method with additional distillation approaches: Decoupled Knowledge Distillation (DKD) (Zhao et al., 2022) and Wasserstein Knowledge Distillation (WKD) (Lv et al., 2024). Both DKD and WKD are orthogonal to DHO since their losses can be applied directly to the outputs of $h_{\text{CE}}$. DKD decouples the target class from non-target classes, and WKD computes a kernel matrix within each class. Both approaches require ground-truth labels, restricting their use to the small labeled dataset $\mathcal{D}^{(l)}$.
Despite these limitations, we conducted experiments on ImageNet (Tab.˜14) and on 8 of the 10 datasets (without Food101, Sun397) as shown in Tab.˜15. The results demonstrate that DHO substantially improves the performance of these state-of-the-art KD methods. Notably, WKD+DHO consistently outperforms DHO alone in most settings, demonstrating the extensibility of DHO due to its simplicity.
Table 14: Performance comparison of DKD and WKD with and without DHO on ImageNet. Numbers in parentheses show improvement over the base method.
Method 1-shot 2-shot 4-shot 8-shot 16-shot
DHO 51.8 52.4 52.6 53.3 54.5
DKD 8.9 15.0 20.6 28.2 34.9
DKD+DHO 47.1 (+38.2) 44.3 (+29.3) 40.9 (+20.3) 40.2 (+12.0) 42.4 (+7.5)
WKD 11.0 17.0 17.0 27.7 34.6
WKD+DHO 53.2 (+42.2) 53.3 (+36.3) 53.3 (+36.3) 54.0 (+26.3) 54.8 (+20.2)
Table 15: Performance comparison of DKD and WKD with and without DHO on 8 datasets (average). Numbers in parentheses show improvement over the base method.
Method 1-shot 2-shot 4-shot 8-shot 16-shot
DHO 60.1 63.7 70.2 75.4 79.1
DKD 28.7 41.3 55.4 66.3 73.1
DKD+DHO 46.2 (+17.5) 51.2 (+9.9) 61.8 (+6.4) 69.8 (+3.5) 74.9 (+1.8)
WKD 30.0 38.2 53.3 66.1 73.6
WKD+DHO 59.6 (+29.6) 64.5 (+26.3) 71.2 (+17.9) 75.7 (+9.6) 79.6 (+6.0)
E.4 Additional Results with Adaptive Weighting
To further explore the potential of DHO, we implement an entropy-based adaptive weighting mechanism. Let the entropy of a probability vector $p\in\Delta^{C-1}$ be $H(p)=-\sum_{c=1}^{C}p_{c}\log p_{c}$. We compute the adaptive weight $\alpha$ as:
$\alpha=\frac{\exp(-H(\hat{p}_{\text{CE}}))}{\exp(-H(\hat{p}_{\text{CE}}))+\exp(-H(\hat{p}_{\text{KD}}))}$ (42)
where $\hat{p}_{\text{CE}}$ and $\hat{p}_{\text{KD}}$ are the output probability vectors from $h_{\text{CE}}$ and $h_{\text{KD}}$, respectively. The final prediction is then computed as $\hat{p}=\alpha\cdot\hat{p}_{\text{CE}}+(1-\alpha)\cdot\hat{p}_{\text{KD}}$.
The intuition behind this approach is that lower entropy (higher confidence) predictions should receive higher weights in the final ensemble. When one head produces more confident predictions than the other, the adaptive weighting mechanism automatically emphasizes the more certain prediction. The results show that the entropy-based adaptive weighting method is not proved to be effective in these experiments. While the adaptive weighting does not consistently outperform the fixed interpolation approach, this is likely due to modern neural networks producing overconfident predictions, making entropy an unreliable proxy for uncertainty without proper calibration. However, we believe adaptive weighting could be beneficial with well-calibrated models or alternative uncertainty measures.
Table 16: Performance of ViT-B/16 and ViT-L/14 distilled from ViT-H/14 with entropy adaptive weighting under different percentages of labeled data.
Method ViT-B/16 ViT-L/14
1% 10% 1% 10%
DHO 81.66 82.78 84.59 85.94
DHO+Ent 81.66 82.65 84.60 85.92
E.5 Out-of-Distribution Evaluation upon fully trained model
We provide the evaluation on out-of-distribution datasets with fully trained model in Tab.˜17. DHO significantly outperformed zero-shot baselines on similar-distribution variants (ImageNet-V2, ImageNet-Sketch) across both ViT-B/16 and ViT-L/14 architectures, but showed performance degradation on out-of-distribution datasets (ImageNet-R, ImageNet-A), suggesting increased distribution overfitting from full model training. Interestingly, ViT-B/16 models distilled from ViT-L/14 handled shifted distributions better than those taught by the larger ViT-H/14 DFN (Fang et al., 2023), despite the latter’s superior performance on shifted distributions such as ImageNet-R and ImageNet-A. We attribute this to the shared training background between ViT-B/16 and ViT-L/14 in the CLIP framework (Radford et al., 2021), which appears to better preserve generalization capabilities during the adaptation. This points to an important insight: our method works best on out-of-distributions when the teacher and student models share similar training distributions, suggesting that successful knowledge distillation also depends on the alignment between teacher and student than just the teacher’s raw capabilities.
Table 17: Accuracy(%) of DHO with full training model on the ImageNet distribution-shifted variants.
Student Model Params (M) Labeled Data Teacher Model Val V2 Sketch R A
ViT-B/16 Student
ViT-B/16 (Radford et al., 2021) 86M zero-shot - 66.7 60.8 46.2 74.0 47.0
ViT-B/16 86M 1% ViT-L/14 78.7 70.1 48.0 70.9 41.1
ViT-B/16 86M 10% ViT-L/14 80.8 71.3 47.4 71.7 41.4
ViT-B/16 86M 1% ViT-H/14 81.6 72.6 50.6 65.5 35.6
ViT-B/16 86M 10% ViT-H/14 82.8 73.6 50.7 67.7 37.8
ViT-L/14 Student
ViT-L/14 (Radford et al., 2021) 304M zero-shot - 75.3 68.3 59.2 86.5 74.6
ViT-L/14 304M 1% ViT-H/14 84.6 77.0 61.5 79.9 60.8
ViT-L/14 304M 10% ViT-H/14 85.9 77.8 61.7 82.8 64.4
Zero-shot VLM
ViT-H (Fang et al., 2023) 632M zero-shot - 83.6 77.2 71.7 92.3 77.4
Appendix F Additional Analysis
F.1 Additional Analysis on Non-Linear Head Design
To further investigate the architectural advantages of dual head optimization, we conducted experiments with non-linear head designs, replacing the linear heads used in our main experiments. We design a non-linear classifier with a sequence of layers: an initial linear projection layer, followed by layer normalization (Ba, 2016), GELU activation (Hendrycks & Gimpel, 2016), dropout (Srivastava et al., 2014), and a final linear classification layer. We compared DHO with three non-linear configurations; DHO+NL-Head-CE: non-linear CE head, DHO+NL-Head-KD: non-linear KD head, and DHO+NL-Head-CE+KD: non-linear both CE and KD heads. All experiments followed the few-shot semi-supervised setting detailed in Section˜C.2.
Table 18: Results of different non-linear head configurations on 11 datasets including ImageNet under few-shot semi-supervision using ResNet-18 with zero-shot teacher (Radford et al., 2021). We report averaged accuracy for 10 visual recognition datasets except the ImageNet.
ImageNet Average of 10 tasks
Configuration 1-shot 2-shot 4-shot 8-shot 16-shot 1-shot 2-shot 4-shot 8-shot 16-shot
DHO (base) 61.7 62.2 62.6 63.8 65.1 58.9 62.2 68.4 73.1 76.5
DHO+NL-Head-CE 61.7 61.9 62.2 63.1 64.8 59.3 62.1 68.0 72.7 76.2
DHO+NL-Head-KD 62.1 62.6 62.9 64.0 65.9 58.3 62.1 67.8 72.4 76.5
DHO+NL-Head-CE+KD 62.0 62.3 62.6 63.8 65.4 59.1 62.2 68.6 72.4 76.5
Figure 12: Results of different non-linear head configurations on 11 datasets including ImageNet under few-shot semi-supervision using ResNet-18 with zero-shot teacher (Radford et al., 2021).
Performance Analysis.
Our experiments revealed key findings regarding head architecture (Tab.˜18, Fig.˜12). On ImageNet, non-linear KD heads consistently outperformed linear ones, suggesting complex architectures better capture teacher predictions. Conversely, non-linear CE heads degraded performance, likely due to overfitting on limited labeled data. While dual non-linear heads outperformed fully linear configurations, they were less effective than non-linearity in the KD head alone.
On the other 10 datasets, optimal configurations varied considerably with no consistently superior approach. This highlights that non-linear transformation effectiveness depends strongly on dataset characteristics and head functionality. Given comparable performance but superior computational efficiency, we adopted linear head architectures for all subsequent experiments.
Table 19: Results on dual-heads interpolation strategy of different non-linear head configurations on ImageNet under 16-shots semi-supervised setting.
Configuration CE Head KD Head Combined
DHO (base) 60.64 61.55 65.37
DHO+NL-Head-CE 60.18 61.39 64.91
DHO+NL-Head-KD 60.95 61.76 65.97
DHO+NL-Head-CE+KD 61.66 61.81 65.59
Head Decomposition Analysis.
Analysis under the 16-shot semi-supervised setting revealed complex relationships between architectural choices and head-wise performance as shown in Tab.˜19. Non-linear CE branches decreased CE head performance from $60.64\%$ to $60.18\%$ despite increased parameters. Conversely, non-linear KD heads improved both heads: CE accuracy increased to $60.95\%$ and KD prediction to $61.76\%$. However, dual non-linear heads reduced combined performance from $65.97\%$ to $65.59\%$, suggesting head specialization may compromise joint feature representation. These findings highlight the complex interplay between architectural decisions and multi-head learning dynamics.
F.2 Additional Dual-Head Analysis
Figure 13: Analysis of DHO on the ImageNet under 16-shot semi-supervised setting. (Left) Maximum probability distributions for predictions from CE head, KD head, and their combined output. (Middle) Prediction agreement diagram analysis, categorizing cases where both heads are correct, only one head is correct, and both heads are incorrect. (Right) Error reduction analysis comparing single-head failure cases against improvements achieved through combined predictions.
In this section, we further analyze the prediction behavior of DHO. As shown in Fig.˜13 (left), despite sharing feature representations, the CE head ($h_{\text{CE}}$), trained on labeled data, produces sharper predictions, whereas the KD head ($h_{\text{KD}}$), guided by teacher distillation, generates smoother distributions. Prediction agreement analysis (Fig.˜13, middle) shows that the two heads agree in 76.2% of cases while complementing each other: the CE head correctly classifies 11.5% of cases where the KD head fails, and vice versa for 12.4%. Error reduction analysis (Fig.˜13, right) further demonstrates that our combined approach reduces failure rates from 11.5% to 2.5% for the KD head and from 12.4% to 5.2% for the CE head, confirming the effectiveness of DHO.
We also present additional qualitative results of DHO, both on ImageNet (see Figs.˜14 and 15) and on other 10 datasets (see Figs.˜16 and 17).
Figure 14: Additional qualitative results on ImageNet under 16-shot semi-supervised setting.
Figure 15: Additional qualitative results on ImageNet under 16-shot semi-supervised setting.
Figure 16: Additional qualitative results on other 10 datasets for models trained under 16-shot semi-supervised setting.
Figure 17: Additional qualitative results on other 10 datasets for models trained under 16-shot semi-supervised setting.
Appendix G The Use of LLMs
We used LLMs solely for light editing such as correcting grammatical errors and polishing some words. They did not contribute to research ideation, experiments, analysis, or substantive writing.